OpenStudy (hang254):

Find an equation for the nth term of the arithmetic sequence. a19 = -92, a20 = 6

OpenStudy (hang254):

@Nurali

OpenStudy (mathstudent55):

What is the common difference?

OpenStudy (irishboy123):

arithmetic series goes : a + (a+d) + (a + 2d) + ...... ***the difference between consecutive terms is d. ***the nth term is a + (n-1), so the 20th term is a + 19 d you now have everything you need to nail this.

OpenStudy (hang254):

86

OpenStudy (mathstudent55):

\(a_1 = a_1~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~ = a_1 + 0d\) \(a_2 = a_1 + 1d~~~~~~~~~~~~~~~~~~~~~~~~~~ = a_1 + 1d\) \(a_3 = a_2 + d = a_1 + d + d ~~~~~= a_1 + 2d\) \(a_4 = a_2 + d = a_1 + 2d + d ~~~= a_1 + 3d\) This suggests an nth term: \(a_n = a_1 + (n - 1)d\)

OpenStudy (mathstudent55):

You have two consecutive terms, a19 and a20. If you subtract a20 - a19, you will find d. What is d?

OpenStudy (mathstudent55):

@IrishBoy123 You meant: "***the nth term is a + (n-1)d, so the 20th term is a + 19 d" ^

OpenStudy (irishboy123):

@mathstudent55 you are eagle eyed!

OpenStudy (hang254):

d = 1 \[a _{n}=-1856+98(n-1)\]

OpenStudy (hang254):

@mathstudent55

OpenStudy (mathstudent55):

a20 - a19 = 98 d = 98 Now we know the common difference is 98.

OpenStudy (mathstudent55):

\(a_{19} = a_1 + 18d\) \(a_19 = a_1 + 18 \times 98\) \(-92 = a_1 + 1764\) \(a_1 = -1856\)

OpenStudy (mathstudent55):

You are correct.

OpenStudy (hang254):

Thank You !

OpenStudy (mathstudent55):

You're welcome.